Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate are as follows:

Standard
Hours
Standard Rate
per Hour
Standard
Cost
24 minutes $5.80 $2.32

During August, 8,420 hours of direct labor time were needed to make 19,800 units of the Jogging Mate. The direct labor cost totaled $47,152 for the month.

Required:

1. What is the standard labor-hours allowed (SH) to makes 19,800 Jogging Mates?

2. What is the standard labor cost allowed (SH × SR) to make 19,800 Jogging Mates?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

5. The budgeted variable manufacturing overhead rate is $4.10 per direct labor-hour. During August, the company incurred $38,732 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month.

(For requirements 3 through 5, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations.)
Answer #1

1)Standars hours(SH) allowed to make 19800 jogging mates

=SH per unit * 19800

=(24/60)*19800

=7920 hours

Note:-24/60 has been taken to convert minutes into hours.

2)Standard Labor Cost (SC) of 19800 jogging mates

=19800*SC per unit

=19800*$2.32

=$45,936

3)Labour Spending Variance

=Standard Cost - Actual Cost(AC)

=$45,936-$47,152

=$1216(U)

4)Labor Rate Variance

=(SR per hour-AR per hour)*AH

=(5.8-5.6)*8420

=$1684(F)

NOTE:- Actual Hours(AH) * Actual Rate per hour(AR)= Actual Cost(AC)

8420 * AR = $47152

AR=47152/8420

AR=$5.6

Labor Efficiency Variance

=(SH-AH)*SR

=(7920-8420)*$5.8

=$2900(U)